# Bridging Model Reference Adaptive Control and Data Informativity

### v2

@@ Abstract
The goal of model reference adaptive control (MRAC) is to ensure that the trajectories of an unknown dynamical system track those of a given reference model. This is done by means of a feedback controller that adaptively changes its gains using data collected online from the closed-loop system. One of the approaches to solve the MRAC problem is to impose conditions on the data that guarantee convergence of the gains to a solution of the so-called matching equations. In the literature, various extensions of the concept of persistent excitation have been proposed in an effort to weaken the conditions on the data required for this convergence. Despite these efforts, it is not well-understood what are the weakest possible data requirements ensuring convergence of MRAC. In this paper, we propose a new framework to study the MRAC problem, using the concept of data informativity. Our main contribution is to provide necessary and sufficient conditions for the asymptotic convergence of the adaptive gains to a solution of the matching equations. These necessary and sufficient conditions can be readily checked online via linear algebra as new data are generated by the closed-loop system. Our results reveal that existing excitation conditions impose stronger requirements on the collected data than required. Notably, the necessary and sufficient conditions provided in this paper are weaker than those for unique system identification.
@@ 1 Introduction
Adaptive control has a rich history, which can be traced back to pioneering work from the 1950s [aseltine1958survey]. Adaptive control aims at dealing with the inevitable uncertainty in control systems by means of adjustable-gain controllers, that are adapted online using the data generated by the closed-loop system. Among the various adaptive control strategies developed over the years, a common approach is model reference adaptive control (MRAC), where the control gains are adjusted so that the closed-loop system tracks a predefined reference model [nguyen2018model].
Meanwhile, offline data-driven control approaches also have a rich history: here, the system to be controlled is uncertain as in adaptive control, but the data are collected offline [willems2012data, de2019formulas]. In this setting, the controller is constructed in one shot using a batch of data. Among the various offline data-driven control strategies, the data informativity approach has recently emerged [van2020data]. Data informativity provides necessary and sufficient conditions under which the data contain sufficient information for control design. In what follows, we provide an overview of representative MRAC approaches and data informativity approaches, so as to clarify the gaps in the respective fields and how these gaps are closed in this work.
@@ 1.1 State of the art on data requirements in MRAC
Despite decades of history, MRAC still remains a prominent research area, with advances continually emerging, such as switched MRAC [di2012hybrid, kersting2017direct, yuan2020lyapunov], robust MRAC [xie2018h, franco2021robust], distributed MRAC [guo2020distributed, yue2023model], partial-state feedback MRAC [song2019partial], and learning-based MRAC [wang2023mrac, annaswamy2023integration], just to mention a few. The key concept common to all MRAC approaches are the matching equations: control gains must exist so that the closed-loop system matches the dynamics of the reference model. Assuming the matching equations have a solution, a pertinent question is whether the online adjustment mechanism converges to such a solution. One of the approaches to MRAC is to impose conditions on the closed-loop data that guarantee convergence of the adjustment mechanism to a solution of the matching equations.
@@
It is well known that the most traditional requirement on the closed-loop data in classical MRAC approaches (direct or indirect [ioannou2006adaptive, duarte1989combined]) goes under the name of persistence of excitation (PE), which requires the excitation provided by the input data to never vanish [boyd1986necessary, annaswamy2021historical]. A very productive line of research has been devoted to relaxing the requirements imposed on data that guarantee convergence to a solution of the matching equations. Representative methods are composite/combined MRAC [lavretsky2009combined], concurrent learning MRAC [chowdhary2010concurrent], and works where PE has been relaxed to rank excitation [chowdhary2013concurrent], finite excitation [cho2017composite, lee2019concurrent] and initial excitation [roy2017combined, katiyar2022initial]. While all these relaxed excitation conditions are sufficient to guarantee convergence to a solution of the matching equations, a natural question is whether they are also necessary. Our work will show that the relaxed excitation conditions in the literature are, in general, not necessary to guarantee convergence to a solution of the matching equations.
@@
It is worth mentioning that although most of the methods above address multi-variable MRAC problems (where the number of inputs and reference signals is larger than one [tao2014multivariable]), such a multi-variable setting requires additional assumptions on the system to arrive at a solution to the problem. It is common in multi-variable MRAC to require a known system’s input matrix [tao2014multivariable, lavretsky2009combined, chowdhary2010concurrent, chowdhary2013concurrent, cho2017composite, lee2019concurrent] or other structural assumptions [costa2003lyapunov, imai2004multivariable, roy2017combined, song2021partial] to handle the uncertainty. Our work will remove such restrictions, requiring no prior knowledge on the input matrix and no other structural assumptions on the unknown system apart from controllability. Moreover, we are able to guarantee convergence to a solution of the matching equations even when the matching equations have a possibly infinite number of solutions, in contrast with existing multi-variable MRAC approaches where the structural assumptions are such that only one solution of the matching equations exists.
@@ 1.2 State of the art on adaptation in data informativity
In the literature on offline data-driven control, a key role is played by Willems et al.’s fundamental lemma [willems2005note]. This result provides a data-dependent representation of the system based on persistently exciting data [de2019formulas, breschi2021direct, schmitz2022willems, pan2022stochastic]. Similar to the quest for relaxed PE requirements in adaptive control, research efforts have been made to relax the requirement for persistently exciting data. An answer in this direction was provided by the data informativity framework [van2020data], demonstrating that PE is not necessary in several data-driven analysis and control problems; also see [van2023informativity] for more details. By adopting the data informativity perspective, [wang2024necessary] studied the requirements that data must have in order to solve an offline model reference control: there, it was shown that persistently exciting data are also not necessary for solving such problem.
However, offline data-driven control is intrinsically different from MRAC, because the data are not generated online by the time-varying closed-loop system obtained from interconnecting the dynamical system and the adaptive feedback law. Indeed, in existing work on the data informativity framework, controllers are constructed offline in one shot, using a batch of data. Online experiment design procedures have been studied in [van2021beyond]: however, such procedures only aim at generating a batch of informative data and do not incorporate a control objective, as is customary in adaptive control. This motivates us to further investigate the relation between MRAC and data informativity. In particular, a key question is: can data informativity be turned from an offline framework to an online one, where data are generated in real-time by the closed-loop system as the controller is adjusted online? In this work, we develop a novel MRAC approach based on the concept of data informativity, thus bridging these two realms. Because data informativity has been shown to give necessary and sufficient conditions for several data-driven control problems, a crucial question is: can necessary and sufficient conditions be obtained for the convergence of MRAC? This work gives a positive answer, thus providing a non-conservative condition on data that allows to guarantee asymptotic convergence to a solution of the matching equations.
@@ 1.3 Contributions
  • We investigate the connection between data informativity for model reference control [wang2024necessary] and data informativity for system identification (cf. Proposition 4.2 & Theorem 4.4). Specifically, we show that data informativity for system identification is sufficient but not necessary for data informativity for model reference control. It becomes necessary only in special cases. This sets the stage for studying MRAC through the lens of data informativity.
  • We propose an MRAC approach that only exploits data informativity for model reference control, without relying on data informativity for system identification or persistence of excitation conditions. The framework relies on data informativity as in work by the authors on offline model reference control [wang2024necessary]. However, we develop new conditions for data informativity for model reference control that can be checked online via linear algebra, as data are generated at every time step by the closed-loop system. This provides an online method for collecting informative data (Lemma 5.7), which is then utilized to design a controller along with an adaptive law. The online collection of informative data marks a key difference with different types of online approaches [rotulo2022online, liu2023online, eising2024data] where the conditions on data are assumed a priori rather than imposed by suitable design of the inputs. We also prove that, with the proposed online method, the controller gains will converge to a solution of the matching equations (Theorem 5.12), and the tracking error is bounded and will converge to zero (Theorem 5.16).
  • The proposed conditions for asymptotic convergence to a solution of the matching equations are necessary and sufficient, and thus do not introduce any conservatism. Furthermore, our MRAC method operates on general multi-variable input-state systems without structural assumptions on the unknown system matrices. Even in scenarios where the ideal controller is not unique, our framework guarantees that the controller parameters will converge to a solution of the matching equations.
@@ 1.4 Outline
The remainder of this paper is structured as follows. In Section II, notation and preliminary results about Willems et al.’s fundamental lemma are provided. Section III presents the mathematical formulation of MRAC. In Section IV, key results on data informativity for system identification and data informativity for model reference control are recalled and compared. The proposed framework for online MRAC is presented in Section V, including necessary and sufficient conditions for the asymptotic convergence to a solution of the matching equations. This section also discusses how the proposed solution relates to the requirements imposed on data in state-of-the-art MRAC methods. In Section VI, numerical examples are provided. Concluding remarks are given in Section VII.
@@ 2.1 Notation
In this section, we recall the fundamental lemma by Willems et al. [willems2005note]. Consider a discrete-time input-state system
@@
Then, for any $t\in\mathbb{N}$, the restricted behavior [^1] Note that the notion of restricted behavior adopted here is slightly different from the literature [willems2005note] because we work with input trajectories of length $t$ and state trajectories of length $t+1$. This is done to facilitate the ensuing discussion. is defined as
@@
([willems2005note, Corollary 2]) Let $t\in\mathbb{N}$ and consider the data $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$. If $(A_{{\rm s}},B_{{\rm s}})$ is controllable and $U_{-}(t)$ is persistently exciting of order $n+1$, then
@@ 3 Problem formulation
where $x_{\rm m}(t)\in\mathbb{R}^{n}$ is the reference state, $r(t)\in\mathbb{R}^{p}$ is the reference input with $p\leq m$, and $A_{\rm m}\in\mathbb{R}^{n\times n},B_{\rm m}\in\mathbb{R}^{n\times p}$ are the reference system matrices. We assume throughout the paper that $(A_{\rm m},B_{\rm m})$ is controllable and $A_{\rm m}$ is Schur. The problem of model reference control is to find matrices $K\in\mathbb{R}^{m\times n}$ and $L\in\mathbb{R}^{m\times p}$ that satisfy the so-called matching equations [ioannou2006adaptive]
$A_{\rm s}+B_{\rm s}K=A_{\rm m},\ B_{\rm s}L=B_{\rm m}.$ (6)
@@
$u(t)=\hat{K}(t)x(t)+\hat{L}(t)r(t),$ (7)
where the gains $\hat{K}(t)\in\mathbb{R}^{m\times n}$ and $\hat{L}(t)\in\mathbb{R}^{m\times p}$ are updated online at each time step using the input-state data generated by (1) with controller (7). Formally, the MRAC problem studied in this work is as follows.
Problem: Given the reference model (5), provide necessary and sufficient conditions under which there exist, for each $t\in\mathbb{Z}_{+}$,
  • integers $i_{t},j_{t}\in\mathbb{N}$, and
  • functions $\alpha_{t}$ and $\beta_{t}$
@@
and the feedback gains
$(\hat{K}(t),\hat{L}(t))=\beta_{t}(U_{-}(t),X(t),\Theta(t))\in\mathbb{R}^{m\times n}\times\mathbb{R}^{m\times p}$ (9)
@@
Here, $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$ is the restricted trajectory of (1) resulting from $x(0)$ and the control input in (7). Moreover, if they exist, provide a recipe to construct $i_{t},j_{t},\alpha_{t}$ and $\beta_{t}$.
@@ 4.1 Conditions on data for system identification and for model reference control
([van2020data]). The data $(U_{-}(t),X(t))$ are informative for system identification if $\Sigma_{(U_{-}(t),X(t))}=\{(A_{\rm s},B_{\rm s})\}$.
@@
([van2020data, Proposition 6]). The data $(U_{-}(t),X(t))$ are informative for system identification if and only if (4) holds.
@@
([wang2024necessary]). The data $(U_{-}(t),X(t))$ are informative for model reference control if there exist control gains $(K,L)$ such that $\Sigma_{(U_{-}(t),X(t))}\subseteq\Sigma^{K,L}$.
@@
We prove that Lemma 3 is a reformulation of [wang2024necessary, Theorem 1]. To show this, note that (11) is equivalent to the existence of matrices $V_{1}\in\mathbb{R}^{T\times n}$ and $V_{2}\in\mathbb{R}^{T\times p}$ such that
$\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}=\begin{bmatrix}X_{-}(t)\\ X_{+}(t)\end{bmatrix}\begin{bmatrix}V_{1}&V_{2}\end{bmatrix}$,
@@
which are precisely the conditions of [wang2024necessary, Theorem 1].
We note that the conditions of Lemma 3 may hold even when the data do not enable the unique identification of $(A_{\rm s},B_{\rm s})$, as shown in [wang2024necessary, Example 1]. This implies that the full rank condition (4) is not always necessary to solve the model reference control problem. In the following, we will show that, if a solution of the matching equations exists, data informativity for system identification is sufficient for data informativity for model reference control but necessary only in some special cases.

### v3

@@ Abstract
The goal of model reference adaptive control (MRAC) is to ensure that the trajectories of an unknown dynamical system track those of a given reference model. This is done by means of a feedback controller that adaptively changes its gains using data collected online from the closed-loop system. One of the approaches to solve the MRAC problem is to impose conditions on the data that guarantee convergence of the gains to a solution of the so-called matching equations. In the literature, various extensions of the concept of persistent excitation have been proposed in an effort to weaken the conditions on the data required for this convergence. Despite these efforts, it is not well-understood what conditions are necessary and sufficient for ensuring convergence of MRAC to a solution of the matching equations. In this paper, we propose a new framework to study the MRAC problem, using the concept of data informativity. Our main contribution is to provide necessary and sufficient conditions for the existence of an adaptive law that guarantees convergence of the gains to a solution of the matching equations, and to provide a recipe for its construction. While existing excitation conditions imply that the system can be uniquely identified from the collected data, our results show that this is not necessary for the convergence of the feedback gains.
@@ 1 Introduction
Adaptive control has a rich history, which can be traced back to pioneering work from the 1950s [3]. Adaptive control aims at dealing with the inevitable uncertainty in control systems by means of adjustable-gain controllers, that are adapted online using the data generated by the closed-loop system. Among the various adaptive control strategies developed over the years, a common approach is model reference adaptive control (MRAC), where the control gains are adjusted so that the closed-loop system tracks a predefined reference model [28].
Meanwhile, offline data-driven control approaches also have a rich history: here, the system to be controlled is uncertain as in adaptive control, but the data are collected offline [42, 11]. In this setting, the controller is constructed in one shot using a batch of data. Among the various offline data-driven control strategies, the data informativity approach has recently emerged [37]. Data informativity provides necessary and sufficient conditions under which the data contain sufficient information for control design. In what follows, we provide an overview of representative MRAC approaches and data informativity approaches, so as to clarify the gaps in the respective fields and how these gaps are closed in this work.
@@ 1.1 State of the art on data requirements in MRAC
Despite decades of history, MRAC still remains a prominent research area, with advances continually emerging, such as switched MRAC [12, 22, 45], robust MRAC [43, 15], distributed MRAC [16, 46], partial-state feedback MRAC [33], and learning-based MRAC [39, 2], just to mention a few. The key concept common to all MRAC approaches are the matching equations: control gains must exist so that the closed-loop system matches the dynamics of the reference model. Assuming the matching equations have a solution, a pertinent question is whether the online adjustment mechanism converges to such a solution. One of the approaches to MRAC is to impose conditions on the closed-loop data that guarantee convergence of the adjustment mechanism to a solution of the matching equations.
@@
It is well known that the most traditional requirement on the closed-loop data in classical MRAC approaches (direct or indirect [19, 13]) goes under the name of persistence of excitation (PE), which requires the excitation provided by the input data to never vanish [5, 1]. In the absence of PE, these classical adaptive schemes can still achieve control objectives like convergence of the tracking error [4], but the adaptive gains may fail to converge to the desired values. In other words, convergence of the gains to a desired solution imposes stronger requirements on data as compared to mere convergence of the tracking error. Motivated by the restrictive nature of the PE, a productive line of research has been devoted to relaxing the requirements imposed on data that guarantee convergence to a solution of the matching equations. Representative methods are composite/combined MRAC [25], concurrent learning MRAC [8], and works where PE has been relaxed to rank excitation [9], finite excitation [7, 26] and initial excitation [31, 21]. While all these relaxed excitation conditions are sufficient to guarantee convergence to a solution of the matching equations, a natural question is whether they are also necessary. Our work will show that such conditions are, in general, not necessary to guarantee convergence to a solution of the matching equations. It is worth mentioning that, while the absence of PE conditions in some adaptive control schemes has been investigated in [24, 23], these results do not apply to the MRAC problem (see the discussion in Section 5.5).
@@
Although most of the methods above address multi-variable MRAC problems (where the number of inputs and reference signals is larger than one [35]), such a multi-variable setting requires additional assumptions on the system to arrive at a solution to the problem. It is common in multi-variable MRAC to require a known system’s input matrix [35, 25, 8, 9, 7, 26] or other structural assumptions [10, 18, 31, 34] to handle the uncertainty. Our work will remove such restrictions, requiring no prior knowledge on the input matrix and no other structural assumptions on the unknown system apart from controllability. Moreover, we are able to guarantee convergence to a solution of the matching equations even when the matching equations have a possibly infinite number of solutions, in contrast with existing multi-variable MRAC approaches where the structural assumptions are such that only one solution of the matching equations exists.
@@ 1.2 State of the art on adaptation in data informativity
In the literature on offline data-driven control, a key role is played by Willems et al.’s fundamental lemma [41]. This result provides a data-dependent representation of the system based on persistently exciting data [11, 6, 32, 29]. Similar to the quest for relaxed PE requirements in adaptive control, research efforts have been made to relax the requirement for persistently exciting data. An answer in this direction was provided by the data informativity framework [37], demonstrating that PE is not necessary in several data-driven analysis and control problems; also see [36] for more details. By adopting the data informativity perspective, [40] studied the requirements that data must have in order to solve an offline model reference control: there, it was shown that persistently exciting data are also not necessary for solving such problem.
However, offline data-driven control is intrinsically different from MRAC, because the data are not generated online by the time-varying closed-loop system obtained from interconnecting the dynamical system and the adaptive feedback law. Indeed, in existing work on the data informativity framework, controllers are constructed offline in one shot, using a batch of data. Online experiment design procedures have been studied in [38]: however, such procedures only aim at generating a batch of informative data and do not incorporate a control objective, as is customary in adaptive control. This motivates us to further investigate the relation between MRAC and data informativity. In particular, a key question is: can data informativity be turned from an offline framework to an online one, where data are generated in real-time by the closed-loop system as the controller is adjusted online? In this work, we develop a novel MRAC approach based on the concept of data informativity, thus bridging these two realms. Because data informativity has been shown to give necessary and sufficient conditions for several data-driven control problems, a crucial question is: can necessary and sufficient conditions be established for the existence of an adaptive law that ensures the convergence of MRAC to a solution of the matching equations? This work gives a positive answer and provides a recipe to design such an adaptive law.
@@ 1.3 Contributions
  • We establish a connection between data informativity for model reference control [40] and data informativity for system identification (Proposition 4.2 & Theorem 4.4). Specifically, we show that data informativity for system identification is sufficient but not necessary for data informativity for model reference control. It becomes necessary only in special cases. This sets the stage for studying MRAC through the lens of data informativity.
  • We provide two equivalent conditions for the existence of an adaptive law guaranteeing asymptotic convergence of the associated gains to a solution of the matching equations (Theorem 5.6). The first condition involves the solvability of the matching equations themselves. The second condition asserts that the underlying system is capable of generating data that are informative for model reference control.
  • Following the existence analysis, we propose an MRAC approach that only exploits data informativity for model reference control, without relying on data informativity for system identification or persistence of excitation conditions. The framework relies on data informativity as in work by the authors on offline model reference control [40]. However, we develop new conditions for data informativity for model reference control that can be checked online via linear algebra, as data are generated at every time step by the closed-loop system. This provides an online method for collecting informative data (Lemma 5.9), which is then utilized to design a controller along with an adaptive law. The online collection of informative data marks a key difference with various online approaches [30, 27, 14] where the conditions on data are assumed a priori rather than imposed by suitable design of the inputs. We also prove that, with the proposed online method, the controller gains will converge to a solution of the matching equations (Theorem 5.14), and the tracking error is bounded and will converge to zero (Theorem 5.18). Notably, our MRAC method operates on general multi-variable input-state systems without structural assumptions on the unknown system matrices. Even in scenarios where the ideal controller is not unique, our framework guarantees that the controller gains will converge to a solution of the matching equations.
@@ 1.4 Outline
The remainder of this paper is structured as follows. In Section II, notation and preliminary results about Willems et al.’s fundamental lemma are provided. Section III presents the mathematical formulation of MRAC. In Section IV, key results on data informativity for system identification and data informativity for model reference control are recalled and compared. The proposed framework for online MRAC is presented in Section V: the framework includes necessary and sufficient conditions for the existence of an adaptive law that guarantees convergence of the gains to a solution of the matching equations, and a recipe for constructing such an adaptive law. This section also discusses how the proposed solution relates to the requirements imposed on data in state-of-the-art MRAC methods. In Section VI, numerical examples are provided. Concluding remarks are given in Section VII.
@@ 2.1 Notation
In this section, we recall the fundamental lemma by Willems et al. [41]. Consider a discrete-time input-state system
@@
Then, for any $t\in\mathbb{N}$, the restricted behavior [^1] Note that the notion of restricted behavior adopted here is slightly different from the literature [41] because we work with input trajectories of length $t$ and state trajectories of length $t+1$. This is done to facilitate the ensuing discussion. is defined as
@@
([41, Corollary 2]) Let $t\in\mathbb{N}$ and consider the data $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$. If $(A_{{\rm s}},B_{{\rm s}})$ is controllable and $U_{-}(t)$ is persistently exciting of order $n+1$, then
@@ 3 Problem formulation
where $x_{\rm m}(t)\in\mathbb{R}^{n}$ is the reference state, $r(t)\in\mathbb{R}^{p}$ is the reference input with $p\leq m$, and $A_{\rm m}\in\mathbb{R}^{n\times n},B_{\rm m}\in\mathbb{R}^{n\times p}$ are the reference system matrices. We assume throughout the paper that $(A_{\rm m},B_{\rm m})$ is controllable and $A_{\rm m}$ is Schur. The problem of model reference control is to find matrices $K\in\mathbb{R}^{m\times n}$ and $L\in\mathbb{R}^{m\times p}$ that satisfy the so-called matching equations [19]
$A_{\rm s}+B_{\rm s}K=A_{\rm m},\quad B_{\rm s}L=B_{\rm m}.$ (6)
@@
$u(t)=\hat{K}(t)x(t)+\hat{L}(t)r(t)+v(t),$ (7)
where the gains $\hat{K}(t)\in\mathbb{R}^{m\times n}$ and $\hat{L}(t)\in\mathbb{R}^{m\times p}$ are updated online at each time step using the input-state data, and the signal $v:\mathbb{Z}_{+}\to\mathbb{R}^{\rm m}$ can be used to excite the system. In this work, we will consider $v$ having finite support, meaning that excitation can only be provided for a finite number of time steps. Formally, the MRAC problem studied in this work is as follows.
Problem 1
Given the reference model (5), provide necessary and sufficient conditions under which there exist, for $t\in\mathbb{Z}_{+}$,
  • integers $i_{t},j_{t}\in\mathbb{N}$,
  • functions $\alpha_{t}$ and $\beta_{t}$, and
  • a finitely supported signal $v$
@@
and the adaptive gains
$\begin{bmatrix}\hat{K}(t)&\hat{L}(t)\end{bmatrix}=\beta_{t}(U_{-}(t),X(t),\Theta(t))\in\mathbb{R}^{m\times n}\times\mathbb{R}^{m\times p}$ (9)
@@
Here, $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$ is the restricted trajectory of (1) resulting from $x(0)$ and the control input in (7). Moreover, if they exist, provide a recipe to construct $i_{t},j_{t},\alpha_{t},\beta_{t}$ and $v$.
@@ 4.1 Conditions on data for system identification and for model reference control
([37]). The data $(U_{-}(t),X(t))$ are informative for system identification if $\Sigma_{(U_{-}(t),X(t))}=\{(A_{\rm s},B_{\rm s})\}$.
@@
([37, Proposition 6]). The data $(U_{-}(t),X(t))$ are informative for system identification if and only if (4) holds.
@@
([40]). The data $(U_{-}(t),X(t))$ are informative for model reference control if there exist control gains $(K,L)$ such that $\Sigma_{(U_{-}(t),X(t))}\subseteq\Sigma^{K,L}$.
@@
We prove that Lemma 3 is a reformulation of [40, Theorem 1]. To show this, note that (11) is equivalent to the existence of matrices $V_{1}\in\mathbb{R}^{t\times n}$ and $V_{2}\in\mathbb{R}^{t\times p}$ such that
$\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}=\begin{bmatrix}X_{-}(t)\\ X_{+}(t)\end{bmatrix}\begin{bmatrix}V_{1}&V_{2}\end{bmatrix}$, (12)
@@
which are precisely the conditions of [40, Theorem 1].
We note that the conditions of Lemma 3 may hold even when the data do not enable the unique identification of $(A_{\rm s},B_{\rm s})$, as shown in [40, Example 1]. This implies that the full rank condition (4) is not always necessary to solve the model reference control problem. In the following, we will show that, if a solution of the matching equations exists, data informativity for system identification is sufficient for data informativity for model reference control but necessary only in some special cases.